OpenStudy (anonymous):

3 times the square of 36y + the square of 100y - 2 times the square of 64y, I got 72 times the square of y. Was I right?

OpenStudy (anonymous):

I'm assuming those are sqaure roots again? You might want to use sqrt instead of square or folks will go down the wrong track.

OpenStudy (saifoo.khan):

can you post a photo your problem here?

OpenStudy (anonymous):

yes square roots, and no I can't post a photo cause it's on a test and it won't let me copy and paste

OpenStudy (anonymous):

\[3\sqrt{36y}+\sqrt{100y}-2\sqrt{64y}\] \[3\sqrt{6^2y}+\sqrt{10^2y}-2\sqrt{8^2y}\] \[18\sqrt{y}+10\sqrt{y}-16\sqrt{y}=12\sqrt{y}\]

OpenStudy (saifoo.khan):

http://take-a-screenshot.org/

OpenStudy (anonymous):

3\[\sqrt{36y}\] +\[\sqrt{100y}\] - 2\[\sqrt{64y}\]

OpenStudy (anonymous):

thanks